# Critical Sections Are Not Per-Thread: A Trace Semantics for Lock-Based Concurrency

### 20 Mar 2026 v2

@@ Abstract.
Locks are a standard mechanism for synchronizing concurrent threads. The standard lock set construction assumes that critical sections are confined to a single thread, and therefore only accounts for locks acquired within that thread. The commonly used notion of a critical section implicitly assumes that protected events belong to the same thread. We show that this assumption is not valid for general C/Pthread executions.
@@ 2. Events and Traces
Traces are well formed and respect the program’s semantics. For example, a lock cannot be acquired twice etc. In the below, we write $l,l^{\prime}$ to refer to lock events and $u,u^{\prime}$ to refer to unlock events. We write $e\in T$ to indicate that $T=[e_{1},...,e_{n}]$ where $e=e_{k}$ for some $k$, defining also that $\mathit{pos}_{T}(e)=k$. For events $e,f\in T$ we define trace order: $e<_{tr}^{T}f$ if $\mathit{pos}_{T}(e)<\mathit{pos}_{T}(f)$.
@@
For all $e=(\upsilon,\mathtt{join}(\tau))\in T$ and all $f=(\tau,op)\in T$ we have $f<_{tr}^{e}$.
@@ 3. Critical Section and Lock Set
We consider a specific program run represented by some well-formed trace $T$. In trace $T$ we wish to identify critical sections that protect events. The entry and exit points of a critical section are identified by matching lock and unlock events $l$ and $u$ where $l$ and $u$ operate on the same lock and are nearest to each other. By well-formedness events $l$ and $u$ must belong to the same thread. In case the program exited prematurely, a lock event $u$ might not have a matching unlock event $l$. In such a case, we consider the last event of the thread the lock event $u$ is in as the exit point.
@@
[Entry and Exit Points] Let $T$ be a well-formed trace with lock event $l=(\tau,\mathtt{lock}(m))\in T$. We refer to $l$ as the entry point of a critical section on lock $m$. The exit point of $l$ in trace $T$ is obtained via function $\mathit{exit}_{T}(l)$ defined below.
@@
For trace $T_{1}$ from Figure 2 we find that $\mathit{exit}_{T_{1}}(e_{6})=e_{11}$, $\mathit{exit}_{T_{1}}(e_{8})=e_{9}$, $\mathit{exit}_{T_{1}}(e_{2})=e_{5}$ and $\mathit{exit}_{T_{1}}(e_{3})=e_{4}$. For trace $T_{3}=[e_{1},e_{6},e_{7},e_{8},e_{9},e_{10},e_{11},e_{2}]$ from Example 2 we find that $\mathit{exit}_{T_{3}}(e_{6})=e_{11}$ and $\mathit{exit}_{T_{3}}(e_{8})=e_{9}$ and $\mathit{exit}_{T_{3}}(e_{2})=e_{2}$.
@@
Let $T$ be a well-formed trace with lock event $l=(t,\mathtt{lock}(m))\in T$. For any reordering $T^{\prime}\in\mathit{crp}(T)$ where $\mathit{exit}_{T}(l)\in T^{\prime}$ we have that $\mathit{exit}_{T}(l)=\mathit{exit}_{T^{\prime}}(l)$.
@@
By construction we have that $\mathit{exit}_{T}(l)=\mathit{exit}_{\mathit{proj}_{T}(\tau)}(l)$. For any $T^{\prime}\in\mathit{crp}(T)$ we have that $\mathit{proj}_{T^{\prime}}(\tau)$ is a prefix of $\mathit{proj}_{T}(\tau)$. By assumption $\mathit{exit}_{T}(l)\in T^{\prime}$ and then we can immediately follow that $\mathit{exit}_{T}(l)=\mathit{exit}_{T^{\prime}}(l)$.
We can argue that critical sections are uniquely identified by their entry and exit points. An event is part of a critical section if the event is enclosed by the entry and its corresponding exit point. In such a situation we say that the event is protected by the lock that is acquired at the entry point.
@@
[Lock Protection] Let $T$ be a well-formed trace with event $e\in T$. We say that $e$ is protected by lock $m$ if there exists lock event $l=(\tau,\mathtt{lock}(m))\in T$ and one of the following cases applies: (LP1) for all $T^{\prime}\in\mathit{crp}(T)$ where $\mathit{exit}_{T}(l)\in T^{\prime}$ we have that $l<_{tr}^{T^{\prime}}e<_{tr}^{T^{\prime}}\mathit{exit}_{T}(l)$, or (LP2) $e=\mathit{exit}_{T}(l)$ and $e\neq(\tau,\mathtt{unlock}(m))$. Case (LP2) covers the case that a lock event $l$ does not have a matching unlock event. Then, $\mathit{exit}_{T}(l)$ equals the last event $\mathit{exit}_{T}(l)$ in $\tau$ and acts as a delimiter for the critical section. In such a case, we consider $\mathit{exit}_{T}(l)$ part of the critical section and protected by lock $m$.
The set of locks that protect an event is commonly referred to as the lock set. Next, we review the existing per-thread definition of a critical section and the resulting lock set construction where the check for enclosed events only considers events that are in the same thread as the exit point. We point out the semantic incompleteness of this assumption and introduce our more general trace-based characterization for which we can establish semantic completeness.
@@ 3.1. Per-Thread Definition
[Per-Thread Critical Section] Let $T$ be a well-formed trace with lock event $l=(\tau,\mathtt{lock}(m))\in T$. We say that $e\in T$ is in the per-thread critical section guarded by $l$ in thread $\tau$, written $e\in CS_{T}^{\tau}(l)$, if one of the following cases applies: (PCS1) $\mathit{proj}_{T}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}\mathbin{\cdot}[\mathit{exit}_{T}(l)]\mathbin{\cdot}T_{3}$ for some traces $T_{1}$, $T_{2}$ and $T_{3}$ where $e\in T_{2}$, or (PCS2) $e=\mathit{exit}_{T}(l)$ and $e\neq(\tau,\mathtt{unlock}(m))$. To deal with ’open’ critical sections where the matching unlock event is missing, we use two cases similar to the prior Definition 3.1.
@@
By assumption, there exists $l=(\tau,\mathtt{lock}(m))\in T$ such that either (PCS1) $\mathit{proj}_{T}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}\mathbin{\cdot}[\mathit{exit}_{T}(l)]\mathbin{\cdot}T_{3}$ for some traces $T_{1}$, $T_{2}$ and $T_{3}$ where $e\in T_{2}$, or (PCS2) $e=\mathit{exit}_{T}(l)$ and $e\neq(\tau,\mathtt{unlock}(m))$.
Consider $T^{\prime}\in\mathit{crp}(T)$ where $\mathit{exit}_{T}(l)\in T^{\prime}$. Each $T^{\prime}\in\mathit{crp}(T)$ maintains the program order condition CRP-PO. Hence, $\mathit{proj}_{T^{\prime}}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}[\mathit{exit}_{T}(l)]\mathbin{\cdot}T_{3}$. Immediately, we can conclude that case (PCS1) implies (LP1) and (PCS2) implies (LP2). Thus, $e$ is protected by lock $m$.
@@
Consider trace $T_{1}$ in Figure 2 where $e_{11}=\mathit{exit}_{T_{1}}(e_{6})$. Due to the fork-join dependency we find that for all reorderings $T^{\prime}\in\mathit{crp}(T_{1})$ where $e_{11}\in T^{\prime}$ we have that event $e_{8}$ is surrounded by events $e_{6}$ and $e_{11}$. That is, $e_{6}<_{tr}^{T^{\prime}}e_{8}<_{tr}^{T^{\prime}}e_{11}$. This means that $e_{8}$ is protected by lock $m$, however, $m\not\in LH_{T_{1}}^{e_{8}}($).
@@ 3.2. Trace-Based Characterization
To guarantee that the lock set is semantically complete, we give a trace-based characterization that lifts the per-thread restriction. For this purpose, we introduce the following ordering relation. We say event $f$ must be preceded by event $e$ under all trace reorderings, written $e<_{crp}^{T}f$, if $\forall T^{\prime}\in\mathit{crp}(T)$ where $f\in T^{\prime}$ we have that $e\in T^{\prime}$ and $e<_{tr}^{T^{\prime}}f$.
@@
[Critical Section] Let $T$ be a well-formed trace with lock event $l=(\tau,\mathtt{lock}(m))\in T$. We say that $e\in T$ is in the critical section guarded by $l$, written $e\in CS_{T}(l)$, if one of the following two cases applies: (CS1) $l<_{crp}^{T}e<_{crp}^{T}\mathit{exit}_{T}(l)$, or (CS2) $e=\mathit{exit}_{T}(l)$ and $e\neq(\tau,\mathtt{unlock}(m))$.
The definition of a critical section is no longer restricted to a single thread. A critical section may cover multiple threads as can be seen by the example in Figure 2.
@@
Let $T$ be a well-formed trace with some event $e=(\tau,op)\in T$. Then, $m\in LH_{T}(\tau){e}$ iff $e$ is protected by lock $m$.
@@
We observe that case (LP1) in Definition 3.1 is equivalent to case (CS1) in Definition 3.2. The same applies to (LP2) and (CS2). Hence, $m\in LH_{T}^{\tau}(e)$ iff $e$ is protected by lock $m$.

### 07 Apr 2026 v4

@@ Abstract.
Locks are a standard mechanism for synchronizing concurrent threads. The standard lock set construction assumes that critical sections are confined to a single thread, and therefore only accounts for locks acquired within that thread. Traditional definitions of critical sections implicitly assume that protected events belong to the same thread. We demonstrate that this assumption does not hold for general C/Pthread executions.
@@ 2. Events and Traces
Traces are well formed and respect the program’s semantics. For example, a lock cannot be acquired twice etc. In the below, we write $l,l^{\prime}$ to refer to lock events and $u,u^{\prime}$ to refer to unlock events. We write $e\in T$ to indicate that $T=[e_{1},...,e_{n}]$ where $e=e_{k}$ for some $k$, defining also that $\mathit{pos}_{T}(e)=k$. For events $e,f\in T$ we define trace order: $e<_{tr}^{T}f$ if $\mathit{pos}_{T}(e)<\mathit{pos}_{T}(f)$. We write $e\leq_{tr}^{T}f$ if $e<_{tr}^{T}f$ or $e=f$.
@@
For all $e=(\upsilon,\mathtt{join}(\tau))\in T$ and all $f=(\tau,op)\in T$ we have $f<_{tr}^{T}e$.
@@ 3. Critical Section and Lock Set
We consider a specific program run represented by some well-formed trace $T$. In trace $T$ we wish to identify critical sections that protect events. The entry and exit points of a critical section are identified by matching lock and unlock events $l$ and $u$ where $l$ and $u$ operate on the same lock and are nearest to each other. By well-formedness events $l$ and $u$ must belong to the same thread. Hence, pairs of matching lock and unlock events can be easily computed via a linear scan of the trace and we write $\mathit{exit}_{T}(l)$ to denote the matching unlock event for lock event $l$. A critical section might be ’open’ in the sense that a lock event $u$ might not have a matching unlock event $l$. In such cases we assume that $\mathit{exit}_{T}(l)$ equals $\bot$. The definition below formally defines $\mathit{exit}_{T}(l)$.
@@
[Entry and Exit Points of Critical Sections] Let $T$ be a well-formed trace with lock event $l=(\tau,\mathtt{lock}(m))\in T$. We refer to $l$ as the entry point of a critical section on lock $m$. The exit point of $l$ in trace $T$ is obtained via the function $\mathit{exit}_{T}(l)$ defined below.
@@
For trace $T_{1}$ from Figure 2 we find that $\mathit{exit}_{T_{1}}(e_{6})=e_{11}$, $\mathit{exit}_{T_{1}}(e_{8})=e_{9}$, $\mathit{exit}_{T_{1}}(e_{2})=e_{5}$ and $\mathit{exit}_{T_{1}}(e_{3})=e_{4}$. For trace $T_{3}=[e_{1},e_{6},e_{7},e_{8},e_{9},e_{10},e_{11},e_{2}]$ from Example 2 we find that $\mathit{exit}_{T_{3}}(e_{6})=e_{11}$ and $\mathit{exit}_{T_{3}}(e_{8})=e_{9}$ and $\mathit{exit}_{T_{3}}(e_{2})=\bot$. The last case shows that a lock event, here $e_{2}$, might lack a matching unlock event.
@@
Let $T$ be a well-formed trace with lock event $l=(t,\mathtt{lock}(m))\in T$. If $\mathit{exit}_{T}(l)\neq\bot$ then for any reordering $T^{\prime}\in\mathit{crp}(T)$ where $\mathit{exit}_{T}(l)\in T^{\prime}$ we have that $\mathit{exit}_{T}(l)=\mathit{exit}_{T^{\prime}}(l)$. If $\mathit{exit}_{T}(l)=\bot$ then for any reordering $T^{\prime}\in\mathit{crp}(T)$ we have that $\mathit{exit}_{T^{\prime}}(l)=\bot$.
@@
By construction we find that (1) $\mathit{exit}_{T}(l)=\mathit{exit}_{\mathit{proj}_{T}(\tau)}(l)$ and (2) for any $T^{\prime}\in\mathit{crp}(T)$ we have that $\mathit{proj}_{T^{\prime}}(\tau)$ is a prefix of $\mathit{proj}_{T}(\tau)$. Hence, if $\mathit{exit}_{T}(l)=\bot$ we immediately find that $\mathit{exit}_{T^{\prime}}(l)=\bot$ for any $T^{\prime}\in\mathit{crp}(T)$.
Consider the case that $\mathit{exit}_{T}(l)\neq\bot$. Consider $T^{\prime}\in\mathit{crp}(T)$ where $\mathit{exit}_{T}(l)\in T^{\prime}$. Then, from (1) and (2) we can immediately follow that $\mathit{exit}_{T}(l)=\mathit{exit}_{T^{\prime}}(l)$.
We can thus argue that critical sections are uniquely identified by their entry and exit points. An event is part of a critical section if the event is enclosed by the entry and its corresponding exit point (if defined). In such a situation we say that the event is protected by the lock that is acquired at the entry point. The set of locks that protect an event is commonly referred to as the lock set. The important question is how to check if an event is ’enclosed’ by another event.
Next, we review the existing per-thread definition of a critical section and the resulting lock set construction where the check for enclosed events only considers events that are in the same thread as the exit point. We point out the semantic incompleteness of this assumption and introduce a more general trace-based characterization that omits the per-thread restriction and is shown to be semantically complete. To establish correctness and completeness, we employ the following characterization of lock protection that does not refer to critical sections explicitly.
@@
[Lock Protection] Let $T$ be a well-formed trace with event $e\in T$. We say that $e$ is protected by lock $m$ if there exists lock event $l=(\tau,\mathtt{lock}(m))\in T$ such that the following two conditions are satisfied:
LP1:
For all $T^{\prime}\in\mathit{crp}(T)$ where $l,e\in T^{\prime}$ we have that $l<_{tr}^{T^{\prime}}e$.
LP2:
There is no unlock event $u=(\tau,\mathtt{unlock}(m))\in T$ such that for some $T^{\prime}\in\mathit{crp}(T)$ where $l,u,e\in T^{\prime}$ we have that $l<_{tr}^{T^{\prime}}u<_{tr}^{T^{\prime}}e$.
Case LP1 guarantees that under all schedules the event $e$ is preceded by the event that acquires lock $m$. Case LP2 guarantees that the lock $m$ is not released in any of these schedules.
@@ 3.1. Per-Thread Definition
[Per-Thread Critical Section] Let $T$ be a well-formed trace with lock event $l=(\tau,\mathtt{lock}(m))\in T$. We say that $e\in T$ is in the per-thread critical section guarded by $l$ in thread $\tau$, written $e\in CS_{T}^{\tau}(l)$, if one of the following two conditions apply:
OPEN-PT:
$\mathit{exit}_{T}(l)=\bot$ and $\mathit{proj}_{T}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}$ for some traces $T_{1}$ and $T_{2}$ where $e\in T_{2}$.
CLOSED-PT:
$\mathit{exit}_{T}(l)\neq\bot$ and $\mathit{proj}_{T}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}\mathbin{\cdot}[\mathit{exit}_{T}(l)]\mathbin{\cdot}T_{3}$ for some traces $T_{1}$, $T_{2}$ and $T_{3}$ where $e\in T_{2}$.
The above states that an event $e$ is in the per-thread critical section guarded by $l$ if $e$ and $l$ belong to the same thread and $e$ occurs after $l$ and if $\mathit{exit}_{T}(l)$ is defined, $e$ occurs before $\mathit{exit}_{T}(l)$ in the trace. This means that the check if $e$ is enclosed by some critical section guarded by $l$ only considers events that are in the same thread as $l$.
@@
By assumption, there exists $l=(\tau,\mathtt{lock}(m))\in T$ and $e\in CS_{T}^{\tau}(l)$.
Case $\mathit{exit}_{T}(l)\neq\bot$: Then, (1) $\mathit{proj}_{T}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}\mathbin{\cdot}[\mathit{exit}_{T}(l)]\mathbin{\cdot}T_{3}$ for some traces $T_{1}$, $T_{2}$ and $T_{3}$ where $e\in T_{2}$.
Consider $T^{\prime}\in\mathit{crp}(T)$ where $l,e\in T^{\prime}$. Each $T^{\prime}\in\mathit{crp}(T)$ maintains the program order condition CRP-PO. Hence, from (1) we derive that $l<_{tr}^{T^{\prime}}e$. This establishes condition LP1.
We establish the condition LP2 by contradiction. Suppose there exists some unlock event $u=(\tau,\mathtt{unlock}(m))\in T$ such that for some $T^{\prime}\in\mathit{crp}(T)$ where $l,u,e\in T^{\prime}$ we have that (2) $l<_{tr}^{T^{\prime}}u<_{tr}^{T^{\prime}}e$. We choose the unlock event $u$ that is nearest to $l$ with no lock event on the same lock in between. The well-formedness assumption guarantees that such an unlock event exists. Hence, $\mathit{exit}_{T^{\prime}}(l)=u$. By Lemma 1 we find that $\mathit{exit}_{T}(l)=u$.
By assumption (1), event $e$ appears before $\mathit{exit}_{T}(l)$ but from (2) we obtain that $e$ appears after $\mathit{exit}_{T}(l)$. Events $e$ and $\mathit{exit}_{T}(l)$ belong to the same thread. Hence, the program order condition CRP-PO is violated and we achieve the desired contradiction.
Case $\mathit{exit}_{T}(l)=\bot$: We find that (3) $\mathit{proj}_{T}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}$ for some traces $T_{1}$ and $T_{2}$ where $e\in T_{2}$.
Consider $T^{\prime}\in\mathit{crp}(T)$ where $l,e\in T^{\prime}$. Each $T^{\prime}\in\mathit{crp}(T)$ maintains the program order condition CRP-PO. Hence, from (3) we derive that $l<_{tr}^{T^{\prime}}e$. This establishes condition LP1.
We establish the condition LP2 by contradiction. Suppose there exists some unlock event $u=(\tau,\mathtt{unlock}(m))\in T$ such that for some $T^{\prime}\in\mathit{crp}(T)$ where $l,u,e\in T^{\prime}$ we have that (2) $l<_{tr}^{T^{\prime}}u<_{tr}^{T^{\prime}}e$. We choose the unlock event $u$ that is nearest to $l$ with no lock event on the same lock in between. The well-formedness assumption guarantees that such an unlock event exists. Hence, $\mathit{exit}_{T^{\prime}}(l)=u$. By Lemma 1 we find that $\mathit{exit}_{T}(l)=u$. This contradicts the assumption that $\mathit{exit}_{T}(l)=\bot$.
For both cases we can establish conditions LP1 and LP2. Thus, we can conclude that $e$ is protected by lock $m$.
@@
Consider trace $T_{1}$ in Figure 2 where $e_{11}=\mathit{exit}_{T_{1}}(e_{6})$. Due to the fork-join dependency we find that for all reorderings $T^{\prime}\in\mathit{crp}(T_{1})$ where $e_{11}\in T^{\prime}$ we have that event $e_{8}$ is always surrounded by events $e_{6}$ and $e_{11}$. That is, $e_{6}<_{tr}^{T^{\prime}}e_{8}<_{tr}^{T^{\prime}}e_{11}$. This means that $e_{8}$ is protected by lock $m$, however, $m\not\in LH_{T_{1}}^{e_{8}}($) because $e_{8}$ arises in a different thread than $e_{6}$.
@@ 3.2. Trace-Based Characterization
To guarantee that the lock set is semantically complete, we give a trace-based characterization that lifts the per-thread restriction when checking if an event is enclosed by another event. For this purpose, we introduce the following ordering relation. Let $e,f$ be two events in trace $T$. We say event $f$ must be preceded by event $e$ under all trace reorderings of $T$, written $e<_{crp}^{T}f$, if $\forall T^{\prime}\in\mathit{crp}(T)$ where $f\in T^{\prime}$ we have that $e\in T^{\prime}$ and $e<_{tr}^{T^{\prime}}f$.
@@
[Critical Section] Let $T$ be a well-formed trace with lock event $l=(\tau,\mathtt{lock}(m))\in T$. We say that $e\in T$ is in the critical section guarded by $l$, written $e\in CS_{T}(l)$, if one of the following two conditions apply:
OPEN:
$\mathit{exit}_{T}(l)=\bot$ and $l<_{crp}^{T}e$.
CLOSED:
$\mathit{exit}_{T}(l)\neq\bot$ and $l<_{crp}^{T}e<_{crp}^{T}\mathit{exit}_{T}(l)$.
The above states that an event $e$ is part of a critical section guarded by $l$ if the event is $e$ is enclosed by the critical section’s entry point $l$ regardless of the schedule of events. If the exit point $\mathit{exit}_{T}(l)$ exit, $e$ is also enclosed by $\mathit{exit}_{T}(l)$. Importantly, event $e$ does not need to arise in the same thread as $l$. Thus, the above definition can deal with critical sections that cover multiple threads as seen by the example in Figure 2.
@@
Let $T$ be a well-formed trace with some event $e=(\tau,op)\in T$. Then, $m\in LH_{T}(e)$ iff $e$ is protected by lock $m$.
@@
$\Rightarrow$”: We consider first the direction from left to right. We have that $m\in LH_{T}(e)$ iff there exists $l=(\tau,\mathtt{lock}(m))\in T$ and $e\in CS_{T}(l)$.
By assumption $l<_{crp}^{T}e$ and thus we can establish condition LP1.
Case $\mathit{exit}_{T}(l)\neq\bot$: We find that (1) $e\leq_{crp}^{T}\mathit{exit}_{T}(l)$. We establish condition LP2 by contradiction via similar reasoning as in the proof of Lemma 2.
Suppose there exists some unlock event $u=(\tau,\mathtt{unlock}(m))\in T$ such that for some $T^{\prime}\in\mathit{crp}(T)$ where $l,u,e\in T^{\prime}$ we have that (2) $l<_{tr}^{T^{\prime}}u<_{tr}^{T^{\prime}}e$. We choose the unlock event $u$ that is nearest to $l$ with no lock event on the same lock in between. The well-formedness assumption guarantees that such an unlock event exists. Hence, $\mathit{exit}_{T^{\prime}}(l)=u$. By Lemma 1 we find that $\mathit{exit}_{T}(l)=u$.
From (1) we derive that the event $e$ appears before $\mathit{exit}_{T}(l)$ but from (2) we obtain that $e$ appears after $\mathit{exit}_{T}(l)$. Events $e$ and $\mathit{exit}_{T}(l)$ belong to the same thread. Hence, the program order condition CRP-PO is violated and we achieve the desired contradiction.
Case $\mathit{exit}_{T}(l)=\bot$: The reasoning that leads to a contradiction is exactly the same as in the proof of Lemma 2.
$\Leftarrow$”: We consider the direction from right to left.
From condition LP1 we can immediately derive that $l<_{crp}^{T}e$. Consider the case that $\mathit{exit}_{T}(l)\neq\bot$. It remains to show that $e<_{crp}^{T}\mathit{exit}_{T}(l)$. Assume the contrary.
This means there exists $T^{\prime}\in\mathit{crp}(T)$ where $\mathit{exit}_{T}(l)<_{tr}^{T^{\prime}}e$. By construction we also find that $l<_{tr}^{T^{\prime}}\mathit{exit}_{T}(l)$. This contradicts condition LP2 and we are done.